OpenStudy (anonymous):

Sarah typed 225 words in five minutes. She then typed 135 words in three minutes. If Sarah types for 24 minutes, how many words does she type? a. What is the answer? b. Explain what the constant of variation is in the problem. Write an equation for the problem. c. What is a constant of variation?

OpenStudy (danjs):

They give you 2 points

OpenStudy (danjs):

(5,225) and (3,135)

OpenStudy (danjs):

You need to figure out the k in y = k*x

OpenStudy (danjs):

The amount of words she types is directly proportional to the time she types

OpenStudy (danjs):

225 = k*5 k = 225 / 5

OpenStudy (danjs):

so the general equation is y = 45*x where x is the time she types, and y is the word count for that time

OpenStudy (danjs):

Anytime you see "constant of variation", it is a graph of a line through the origin (0,0)

OpenStudy (danjs):

a) types for 24 min y = 45 * 24 = 1080 words typed b)k = 45, the constant of variation is the slope of the line through the origin. It also means that for each min she types, she produces 4 words. c) the equation is y = k*x y = 45x

OpenStudy (danjs):

45 words*

OpenStudy (anonymous):

Thank you!

OpenStudy (danjs):

no prob, anytime